# Amorphous silicon structures generated using a moment tensor potential and the activation relaxation technique nouveau

### v1

@@ III Results and discussion
Fig. 2: Coordination numbers of input configurations and ARTn-MTP optimized configurations (final configurations). All non-4-fold coordinated atoms, including under- and over-coordinated atoms, are considered defective. The models analyzed include 216, 512 , 1000 and 4096 atoms.
@@
Fig. 3: Comparison of the local structural properties of input configurations (top panels: a, b, and c) with the final configuration (100% coordinated) (bottom panels: d, e, and f) generated using the ARTn-MTP coupling scheme. The structural properties analyzed include angular distribution, radial distribution, and bond length distribution for a system comprising 216 atoms. Experimental data and data from the 100k-GAP18 model 61 are included to assess the quality of our 100% 4-fold coordinated 216-atom a-Si model
@@
Models of size 216 atoms, 512 atoms, 1000 atoms and 4096 atoms were generated as detailed in table 1. We analyze these configurations with respect to short-range order, medium-range order, electronic properties as well as mechanical properties. Fig. 2 displays the defects in both the input and final configurations. The input configurations exhibit a higher prevalence of coordination defects than the final configurations: approximately 70% for the 216-R1, 216-R2, and 512-R1 models, compared to around 30% for the 512-R2 and 1000-R models, as illustrated by the blue bars. The green bars represent the total number of defects present in the final configurations.
@@
In all ARTn-MTP a-Si models, the final percentage of defective atoms remained below 2%, with the exception of the 1000-R and 4096-R-MD models, which exhibit a final defect percentage of approximately 3%. The initial 4-fold coordination number (shown by the orange bars in Fig. 2) is approximately 30% for the 216-R1, 216-R2, and 512-R1 models, and around 68% for the 512-R2 and 1000-R models. In contrast, the final configurations display markedly higher 4-fold coordination numbers (depicted by the red bars in Fig. 2 and Table 2), with values of 99.07%, 100%, 98.04%, 98.04%, 97.20%, and 97.80% for the 216-R1, 216-R2, 512-R1, 512-R2, 1000-R and 4096-R-MD models, respectively. No atoms with fewer than 3 neighbors or more than 5 neighbors were observed.
@@
To further contextualize the number of coordination defects in the ARTn-MTP a-Si models, we present a comparison in Fig. 5, which includes data from the literature as well. First, all the models presented show no atoms with fewer than 3 neighbors or more than 5 neighbors. Secondly, all of the ARTn-MTP generated models exhibit fewer floating and dangling bonds compared to all other methods, except the WWW model and 100k-GAP18 model. Notably, the 216-R2 model is absent from Fig. 5 because it contains 100% four-fold coordination. To our knowledge, no other a-Si models exhibit 100% four-fold coordination, except for some WWW models generated with a Keating-type potential 78, 23
@@
The last column of Table 2 presents the fraction of atoms in a crystalline environment (cubic or hexagonal diamond). As shown, most of the generated a-Si models, particularly those from MD simulations, contain some proportion of crystallinity. Although WWWW is known to produce large a-Si models without coordination defects, the WWW a-Si models still exhibit traces of crystallinity. We identified four models—216-R1, 512-R1, 512-R2, and 1000-R—that are free from crystallinity. Furthermore, as indicated in Table  3, both the 1000-aSi-SW initial and 1000-aSi-SW final configurations also contain no crystallinity. It is important to note that all four of our a-Si models without crystallinity, including the 1000-aSi-SW initial model, were generated starting from a random state without any prior MD prerun, meaning their boxes were initially filled randomly. From Table 2, we note three potential trends. (i) this fraction does not seem to depend much on the system size. (ii) Starting ARTn from a random state seems to lead to a less crystalline state than starting from a melt-and-quench state. (iii) Fewer coordination defects seem to favor a higher crystallinity content. Now that, with reliable ML potentials, we can generate similar quality models using different pathways (non-dynamic melt-and-quench methods), we can start examining the link between local crystallinity, defects and the ideal continuous-random networks. This can lead to understanding why, for example, amorphous silicon cannot be produced experimentally from the liquid phase and crystallizes well below melting as well as whether the CRN model corresponds really to ideal a-Si. The ML potential 62, 70, which leads to high-quality models, makes further study on this question now possible.
@@
Fig. 5: Proportion of coordination defects in various a-Si models generated by different methods. 216-R1 has no floating bond and 216-R2 is 100 % fully 4-fold coordinated. No atoms with fewer than 3 neighbors or more than 5 neighbors have been recorded in any model.
Fig. 6: FWHM of the bond-angle distribution. The FWHM values are computed by fitting a Gaussian distribution to the bond angle distributions. The red line denotes the experimental value of FWHM 20.
@@ III.1 Short-range order structural analysis of the models
Fig. 3 illustrates the structural properties of the 216-R2 models, including the angular distribution function, radial distribution function, and bond length distribution. Panels a), b), and c) display these properties for the randomly packed input configuration (216-R2), while panels d), e), and f) show the properties of the final configurations (216-R2) produced by ARTn-MTP. The final 216-R2 configuration is 100% fully coordinated, with no coordination defects present. As shown in Fig. 3 a), b), and c), the data for the input configurations are randomly distributed and widespread. In contrast, the data for the final configurations, depicted in Fig. 3 d), e), and f), follow well-defined distributions. The shape of the radial distribution function matches the experimental data, though the peak is somewhat overestimated. This indicates that the Si-Si pair distances in our model are widely distributed around the experimental pair separation distance. The bond angle distribution function is centered around the tetrahedral value of 109.5° with a spread of 9.96°. A Gaussian fit to the data reveals a full width at half-maximum (FWHM) of 23.45°. The half width at half-maximum (HWHM) is 11.73°, which is similar to the experimental HWHM value of 11° 20. For the bond length, the data for the input configuration (3 c) are widespread. In contrast, for the final configuration (3 f), the bond lengths follow a Gaussian distribution centered around the experimental value of 2.35 Å. The mean bond length is 2.36 Å, with a spread of 0.03 Å. The angular distribution function, radial distribution function, and bond length distribution of the final model (216-R2) were compared with those of the 100k-GAP18 configuration, which contains 100,000 atoms. As shown in Fig. 3 (d), (e), and (f), these properties align well with those of the larger 100k-GAP18 model. This demonstrates the high quality of our fully coordinated a-Si model. Beyond the analysis of the fully coordinated configuration, we also explore the short-range order structure in our other models. The detailed results are presented in Fig. 4 and Tab. 2. Once again, the structural properties—including the radial distribution function (Fig. 4 a), bond angle distribution function (Fig. 4 b), and bond length distribution (Fig. 4 c)—closely match experimental data and those of 100k-GAP18 models. The data for all the models presented in Fig. 4 show similar behavior, with the exception of the Si-Si bond length distribution. Notably, the bond lengths for two of our models—216-R1 (in green) and 512-R1 (in red)—are shifted. These two models have a density of 2.20 g/cm³, which is lower than that of the other models, which is 2.28 g/cm³. Since amorphous silicon features local tetrahedral bonding, we quantify the deviation from ideal tetrahedral geometry by calculating the FWHM for all models included in this study, as presented in Fig. 6. The FWHM values for all our models closely align with the experimental data, as well as with the WWW model and the 100k-GAP18 models.
@@ III.2 Medium-range order structural analysis of the models
Fig. 7: Structure factor for ARTn-MTP generated models compared with that obtained from experimental data and 100k-GAP18 atom simulations.
Fig. 8: Ring statistics for ARTn-MTP generated models as well as those derived from the 512-WWW model and 100k-GAP18 atom models.
Ring statistics are illustrated in Fig. 8. Crystalline silicon, characterized by its diamond-type structure, primarily features 6-membered rings, while amorphous silicon tends to favor rings of sizes 5, 6, and 7. Rings with fewer than 5 members or more than 7 are typically regarded as ring defects. Notably, no 3-membered rings are observed in any of the models. Additionally, very few 4-membered rings, which are known to introduce structural stress, were recorded across all models. As shown in Fig. 8 (a-f), each model predominantly contains 5, 6, and 7-membered rings, consistent with existing literature. These ring sizes are energetically favorable for a-Si. Furthermore, all models exhibit a few larger ring defects with sizes greater than 7, as detailed in Fig. 8.
@@ III.3 Refinement of ART-SW optimized model: 1000-aSi-SW
Fig. 9: Local structural properties of ARTn-MTP refined configurations (1000-aSi-SW final) compared to ART-SW optimized configurations (1000-aSi-SW initial). a) Initial and final counts of 3-fold and 5-fold defects, b) bond angle distribution function, c) radial distribution function, and d) structure factor. For figures (b), (c), and (d), experimental data and results from the 100k-GAP-18 atom model are included to assess the quality of both the initial and final configurations of the 1000-aSi-SW model. The ART-SW optimized configurations served as the input for the ARTn-MTP refined configurations.
As we first test the capability of ARTn coupled with MLIP to model amorphous systems, it is essential to explore several scenarios. A key consideration is the starting input configuration (see Table 1). Starting from a completely random state presents a challenge, particularly for larger systems. Ideally, the process should begin from a local minimum, which can be obtained through methods such as MD runs, CG relaxation, or ARTn coupled with a semi-empirical model. In this section, we present preliminary results that demonstrate the capability of ARTn-MTP to push the limits of relaxation for a configuration already optimized by ARTn coupled with the Stillinger-Weber (SW) potential 79.
A previously published a-Si model 22, 38 using ARTn with the SW potential 79, modified by Vink 80, served as input for an ARTn-MTP process. The 1000-atom ARTn-SW a-Si configuration (1000-aSi-SW initial) was refined through ARTn-MTP, resulting in the (1000-aSi-SW final) configuration. The goal was to assess the ARTn-MTP process’s ability to refine configurations optimized with a semi-empirical potential. The energy dropped by -23.25 meV/atom, and structural properties, shown in Fig. 9 and summarized in Tab. 3, reveal significant improvements. The number of dangling and floating bonds decreased, reducing total defects by 3%. Bond angles (Fig. 9 b) are closer to the ideal 109.5°, with a narrower spread in the final model. Bond lengths were concentrated around 2.38 Å, aligning with experimental values, and the radial distribution function in Fig. 9 c confirmed this. The 1000-aSi-SW final and 100k-GAP18 models showed better agreement with experimental data than the initial configuration. The structure factor, in Fig. 9 d, also matched experimental S(q). The excess energy ($\Delta E$) decreased from 0.2021 eV/atom (initial) to 0.1789 eV/atom (final), falling within the experimental range of 0.135 – 0.205 eV/atom 81. The final configuration, with lower excess energy, is more stable and realistic. The bulk modulus increased from 65.8573 GPa (initial) to 68.8690 GPa (final), indicating enhanced stiffness. These results demonstrate that the ARTn-MTP coupling scheme effectively relaxes a semi-empirically optimized amorphous configuration.
@@ III.4 Stiffness
Fig. 10: Bulk modulus for ARTn-MTP generated models as well as models from the literature.
We calculated the bulk modulus $B$ of our models and of the models found in the literature. The results are presented in Fig. 10. We include the bulk modulus value obtained from density functional theory (DFT) simulations reported in Ref. 82, which is 82.5 GPa. For our 216-R2 configuration with 100% 4-fold coordination, we obtained a bulk modulus of 80.01 GPa. The 512-R2 model yields a bulk modulus of 77.53 GPa, closely matching the 512-WWW model’s value of 77.41 GPa. In contrast, other models such as 512-MD, 512-INV, and 512-RMC exhibit lower bulk modulus values of 71.18 GPa, 69.00 GPa, and 70.90 GPa, respectively. The 1000-R model demonstrates a bulk modulus of 76.46 GPa. Our largest models, the 4096-R-MD and 100k-GAP18, yield bulk modulus values of 77.93 GPa and 79.56 GPa, respectively. All of these values were obtained using our unified machine learning potential developed for the Si, O, and SiO2 systems. Other potentials may yield different results. To our knowledge, there are no clear experimental benchmark values for the bulk modulus of amorphous silicon.

### v2

@@ III Results and discussion
Fig. 2: Comparison of the local structural properties of input configurations (top panels: a, b, and c) with the final configuration (100% coordinated) (bottom panels: d, e, and f) generated using the ARTn-MTP coupling scheme. The structural properties analyzed include angular distribution, radial distribution, and bond length distribution for a system comprising 216 atoms. Experimental data and data from the 100k-GAP18 model 67 are included to assess the quality of our 100% 4-fold coordinated 216-atom a-Si model
@@
Models of size 216 atoms, 512 atoms, 1000 atoms and 4096 atoms were generated as detailed in table 1. We analyze these configurations with respect to short-range order, medium-range order, as well as mechanical properties. First, we would like to emphasize that, the input configurations exhibit a higher prevalence of coordination defects than the final configurations: approximately 70% for the 216-R1, 216-R2, and 512-R1 models, compared to around 30% for the 512-R2 and 1000-R models.
@@
In all ARTn-MTP a-Si models, the final percentage of defective atoms remained below 2%, with the exception of the 1000-R and 4096-R-MD models, which exhibit a final defect percentage of approximately 3%. The initial 4-fold coordination number is approximately 30% for the 216-R1, 216-R2, and 512-R1 models, and around 68% for the 512-R2 and 1000-R models. In contrast, the final configurations display markedly higher 4-fold coordination numbers (Table 2), with values of 99.07%, 100%, 98.04%, 98.04%, 97.20%, and 98.05% for the 216-R1, 216-R2, 512-R1, 512-R2, 1000-R and 4096-R-MD models, respectively. No atoms with fewer than 3 neighbors or more than 5 neighbors were observed.
@@
To further contextualize the number of coordination defects in the ARTn-MTP a-Si models, we present a comparison in Table 2, which includes data from MD simulation and the literature as well. First, all the models presented show no atoms with fewer than 3 neighbors or more than 5 neighbors. Secondly, all of the ARTn-MTP generated models exhibit fewer floating and dangling bonds compared to all other methods, except the WWW model and 100k-GAP18 model. Notably, the 216-R2 model exhibit 100% four-fold coordination. To our knowledge, no other a-Si models exhibit 100% four-fold coordination, except for some WWW models generated with a Keating-type potential 85, 29
@@
The last column of Table 2 presents the fraction of atoms in a crystalline environment (cubic or hexagonal diamond). As shown, most of the generated a-Si models, particularly those from MD simulations, contain some proportion of crystallinity. Although WWW is known to produce large a-Si models without coordination defects, the WWW a-Si models still exhibit traces of crystallinity. We identified four models—216-R1, 512-R1, 512-R2, and 1000-R—that are free from crystallinity. Furthermore, as indicated in Table  3, both the 1000-aSi-SW initial and 1000-aSi-SW final configurations also contain no crystallinity. It is important to note that all four of our a-Si models without crystallinity, including the 1000-aSi-SW initial model, were generated starting from a random state without any prior MD prerun, meaning their boxes were initially filled randomly. From Table 2, we note three potential trends. (i) this fraction does not seem to depend much on the system size. (ii) Starting ARTn from a random state seems to lead to a less crystalline state than starting from a melt-and-quench state. (iii) Fewer coordination defects seem to favor a higher crystallinity content. Now that, with reliable ML potentials, we can generate similar quality models using different pathways (non-dynamic melt-and-quench methods), we can start examining the link between local crystallinity, defects and the ideal continuous-random networks. This can lead to understanding why, for example, amorphous silicon cannot be produced experimentally from the liquid phase and crystallizes well below melting as well as whether the CRN model corresponds really to ideal a-Si. The ML potential 68, 76, which leads to high-quality models, makes further study on this question now possible.
@@
Fig. 4: FWHM of the bond-angle distribution. The FWHM values are computed by fitting a Gaussian distribution to the bond angle distributions. The bars are hatched for our purely MD-generated a-Si models. The red line denotes the experimental value of FWHM 20.
@@ III.1 Short-range order structural analysis of the models
Fig. 2 illustrates the structural properties of the 216-R2 models, including the angular distribution function, radial distribution function, and bond length distribution. Panels a, b, and c display these properties for the randomly packed input configuration (216-R2), while panels d, e, and f show the properties of the final configurations (216-R2) produced by ARTn-MTP. The final 216-R2 configuration is 100% fully coordinated, with no coordination defects present. As shown in Fig. 2 a, b, and c, the data for the input configurations are randomly distributed and widespread. In contrast, the data for the final configurations, depicted in Fig. 2 d, e, and f, follow well-defined distributions. The shape of the radial distribution function matches the experimental data, though the peak is somewhat overestimated. This indicates that the Si-Si pair distances in our model are widely distributed around the experimental pair separation distance. The bond angle distribution function is centered around the tetrahedral value of 109.5° with a spread of 9.96°. A Gaussian fit to the data reveals a full width at half-maximum (FWHM) of 23.45°. The half width at half-maximum (HWHM) is 11.73°, which is similar to the experimental HWHM value of 11° 20. For the bond length, the data for the input configuration (Fig. 2 c are widespread. In contrast, for the final configuration (Fig. 2 f, the bond lengths follow a Gaussian distribution centered around the experimental value of 2.35 Å. The mean bond length is 2.36 Å, with a spread of 0.03 Å. The angular distribution function, radial distribution function, and bond length distribution of the final model (216-R2) were compared with those of the 100k-GAP18 configuration, which contains 100,000 atoms. As shown in Fig. 2 d, e, and f, these properties align well with those of the larger 100k-GAP18 model. This demonstrates the high quality of our fully coordinated a-Si model. Beyond the analysis of the fully coordinated configuration, we also explore the short-range order structure in our other models. The detailed results are presented in Fig. 3 and Tab. 2. Once again, the structural properties—including the radial distribution function (Fig. 3 a), bond angle distribution function (Fig. 3 b), and bond length distribution (Fig. 3 c)—closely match experimental data and those of 100k-GAP18 models. The data for all the models presented in Fig. 3 show similar behavior, with the exception of the Si-Si bond length distribution. Notably, the bond lengths for two of our models—216-R1 (in green) and 512-R1 (in red)—are shifted. These two models have a density of 2.20 g/cm³, which is lower than that of the other models, which is 2.28 g/cm³. Since amorphous silicon features local tetrahedral bonding, we quantify the deviation from ideal tetrahedral geometry by calculating the FWHM for all models included in this study, as presented in Fig. 4. The FWHM values for all our models closely align with the experimental data, as well as with the WWW model and the 100k-GAP18 models.
@@ III.2 Medium-range order structural analysis of the models
Fig. 5: Structure factor for ARTn-MTP generated models compared with that obtained from experimental data from 13 and 100k-GAP18 atom simulations.
Fig. 6: Ring statistics for ARTn-MTP generated models as well as those derived from the 512-WWW model and 100k-GAP18 atom models. The total ring count is normalized by the number of atoms in the model.
Ring statistics are illustrated in Fig. 6. Crystalline silicon, characterized by its diamond-type structure, primarily features 6-membered rings, while amorphous silicon tends to favor rings of sizes 5, 6, and 7. Rings with fewer than 5 members or more than 7 are typically regarded as ring defects. Notably, no 3-membered rings are observed in any of the models. Additionally, very few 4-membered rings, which are known to introduce structural stress, were recorded across all models. As shown in Fig. 6 a-f, each model predominantly contains 5, 6, and 7-membered rings, consistent with existing literature. These ring sizes are energetically favorable for a-Si. Furthermore, all models exhibit a few larger ring defects with sizes greater than 7, as detailed in Fig. 6.
@@ III.3 Refinement of ART-SW optimized model: 1000-aSi-SW
Fig. 7: Local structural properties of ARTn-MTP refined configurations (1000-aSi-SW final) compared to ART-SW optimized configurations (1000-aSi-SW initial). a) Initial and final counts of 3-fold and 5-fold defects, b) bond angle distribution function, c) radial distribution function, and d) structure factor. For figures (b), (c), and (d), experimental data 13 and results from the 100k-GAP-18 atom model are included to assess the quality of both the initial and final configurations of the 1000-aSi-SW model. The ART-SW optimized configurations served as the input for the ARTn-MTP refined configurations.
As we first test the capability of ARTn coupled with MLIP to model amorphous systems, it is essential to explore several scenarios. A key consideration is the starting input configuration (see Table 1). Starting from a completely random state presents a challenge, particularly for larger systems. Ideally, the process should begin from a local minimum, which can be obtained through methods such as MD runs, CG relaxation, or ARTn coupled with a semi-empirical model. In this section, we present preliminary results that demonstrate the capability of ARTn-MTP to push the limits of relaxation for a configuration already optimized by ARTn coupled with the Stillinger-Weber (SW) potential 86.
A previously published a-Si model 28, 44 using ARTn with the SW potential 86, modified by Vink 87, served as input for an ARTn-MTP process. The 1000-atom ARTn-SW a-Si configuration (1000-aSi-SW initial) was refined through ARTn-MTP, resulting in the (1000-aSi-SW final) configuration. The goal was to assess the ARTn-MTP process’s ability to refine configurations optimized with a semi-empirical potential. The energy dropped by -23.25 meV/atom, and structural properties, shown in Fig. 7 and summarized in Tab. 3, reveal significant improvements. The number of dangling and floating bonds decreased, reducing total defects by 3%. Bond angles (Fig. 7 b) are closer to the ideal 109.5°, with a narrower spread in the final model. Bond lengths were concentrated around 2.38 Å, aligning with experimental values, and the radial distribution function in Fig. 7 c confirmed this. The 1000-aSi-SW final and 100k-GAP18 models showed better agreement with experimental data than the initial configuration. The structure factor, in Fig. 7 d, also matched experimental S(q). The excess energy ($\Delta E$) decreased from 0.2021 eV/atom (initial) to 0.1789 eV/atom (final), falling within the experimental range of 0.135 – 0.205 eV/atom 88. The final configuration, with lower excess energy, is more stable and realistic. The bulk modulus increased from 65.8573 GPa (initial) to 68.8690 GPa (final), indicating enhanced stiffness. These results demonstrate that the ARTn-MTP coupling scheme effectively relaxes a semi-empirically optimized amorphous configuration.
@@ III.4 Stiffness
Fig. 8: Mechanical properties of the models: a) Bulk modulus of ARTn-MTP generated models, MD generated models (hatched bars), and literature models.b) Elastic constants of ARTn-MTP generated a-Si models and MD simulated models (hatched bars).
We calculated the bulk modulus $B$ of our models and of the models found in the literature. The results are presented in Fig. 8 a. We include the bulk modulus value obtained from density functional theory (DFT) simulations reported in Ref. 89, which is 82.5 GPa. The DFT value was computed using the local density approximation (LDA) and hard, norm-conserving pseudopotentials 89. For our 216-R2 configuration with 100% 4-fold coordination, we obtained a bulk modulus of 80.01 GPa. The 512-R2 model yields a bulk modulus of 77.53 GPa, closely matching the 512-WWW model’s value of 77.41 GPa. In contrast, other models such as 512-MD, 512-INV, and 512-RMC exhibit lower bulk modulus values of 71.18 GPa, 69.00 GPa, and 70.90 GPa, respectively. The 1000-R model demonstrates a bulk modulus of 76.46 GPa. Our largest models, the 4096-R-MD and 100k-GAP18, yield bulk modulus values of 77.93 GPa and 79.56 GPa, respectively.For our a-Si models generated solely using MD simulation, the bulk moduli are as follows: 76.54 GPa for 216-MD1, 76.41 GPa for 216-MD2, 75.51 GPa for 512-MD1, 73.77 GPa for 512-MD2, 73.88 GPa for 1000-MD, and 74.66 GPa for 4096-MD. With the exception of 216-MD2, which shows no crystallinity, the other MD models contain significant proportions of crystallinity. The bulk modulus values presented here reflect the quality of our amorphous models. It is important to note that the bulk modulus for amorphous materials depends heavily on the preparation method and the inherent lack of periodicity in the structure. As a result, the value of the bulk modulus is not unique and varies across different models. The average bulk modulus of all the models, including our ARTn models, MD models, and those from the literature, is 75.47 GPa, with a spread of 3.63 GPa. All the bulk modulus values, including the mean, are notably lower than the bulk modulus of crystalline silicon, which is 98.3 GPa 90. A value of 59 GPa for the experimental bulk modulus of hydrogenated amorphous silicon (a-Si:H) has also been reported 91. This value is significantly lower than the bulk modulus values observed in all of the models presented. All these values were obtained using our unified machine learning potential developed for the Si, O, and SiO2 systems. Other potentials may yield different results. To our knowledge, there are no clear experimental benchmark values for the bulk modulus of amorphous silicon. In addition, we also assessed the elastic constants of all our models, and the results are presented in Fig. 8 b. As can be noted, strong dispersion is not noticeable. For our ARTn-MPT models, the mean values (with spreads) are as follows: 118.46 GPa (4.77 GPa) for C11, 53.99 GPa (0.00 GPa) for C12, and 32.05 GPa (0.79 GPa) for C44. For our a-Si models generated solely with MD, the corresponding values are 119.87 GPa (1.45 GPa) for C11, 52.76 GPa (1.08 GPa) for C12, and 33.24 GPa (0.38 GPa) for C44. These values lie within the same range as those reported in previous works: 121 GPa for C11, 43 GPa for C12, 40 GPa for C44, and 69 GPa for Bulk modulus 30. Thus, while mechanical properties are highly dependent on the quality of the potential, we can argue that the computed values presented here fall within the range of the true mechanical properties of real a-Si models.